invested 10000$ for 1 year party at 7% and part 6.25% if she earned a total interest of 656.50$ how much was invested at each rate?
Answer by mananth(16946)

Part I 7.00% per annum ------------- Amount invested =x
Part II 6.25% per annum ------------ Amount invested = y
10000
Interest----- 656.50

Part I 7.00% per annum ---x
Part II 6.25% per annum ---y
Total investment
x + 1 y= 10000 -------------1
Interest on both investments
7.00% x + 6.25% y= 656.5
Multiply by 100
7 x + 6.25 y= 65650.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -70000.00
Add this to (2)
0 x -0.75 y= -4350
divide by -0.75
y = 5800
Part I 7.00% $ 4200
Part II 6.25% $ 5800

CHECK
4200 --------- 7.00% ------- 294.00
5800 ------------- 6.25% ------- 362.50
Total -------------------- 656.50

m.ananth@hotmail.ca